### v2

@@ A.8 Proof of Lemma 3
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=-\frac{\kappa(s,\mathbf{p},\tilde{\mu})(1-q)+q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta)e^{\frac{2\mu s}{\sigma^{2}}}}{q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1}.$
To find the indifferent voter, we need to solve
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}\Leftrightarrow s=s^{+}\equiv\frac{\sigma^{2}\ln\left(1+\frac{\gamma}{q(\beta-\gamma+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta+\Delta))}\right)}{2\mu}.$
Next consider a voter with $s<0$. It follows from Lemma 1 that such a voter would either like to choose $\tilde{\mu}=\infty$ or $\tilde{\mu}^{*}=\mu$. Moreover, by (7), when $\tilde{\mu}=\mu$, utility is the lowest when $s=0$. Hence, if a voter receiving $s=0$ has no incentive to distort $\mu$, then no voter with $s<0$ has an incentive to do so. The payoff of not distorting is
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu\wedge s=0}=-\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta).$
If, however, a voter with $s<0$ chooses to distort beliefs, he receives
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\infty\wedge s<0}=-\kappa(s,\mathbf{p},\tilde{\mu})-\gamma.$
Hence, if $\gamma\geq\gamma^{-}\equiv q(\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu}))$, then no voter with $s<0$ chooses $\tilde{\mu}\neq\mu$. However, if $\gamma<\gamma^{-}$, then there exists a voter with $s=s^{-}<0$ who is indifferent between distorting and not distorting. To find the indifferent voter, we need to solve
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu}=\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\infty\wedge s<0}\Leftrightarrow s=s^{-}\equiv\frac{\sigma^{2}\ln\left(\frac{\gamma(q-1)}{q(\beta(\kappa(s,\mathbf{p},\tilde{\mu})-1)+\gamma-\Delta+\kappa(s,\mathbf{p},\tilde{\mu}))}\right)}{2\mu}.$
Any voter with $s<s^{-}$ chooses $\tilde{\mu}=\mu$, while any voter with $s\in(s^{-},0)$ chooses $\tilde{\mu}=0$.∎
A.9 Proof of Proposition 5
In an inactivity equilibrium, we have $\kappa(s,\mathbf{p},\tilde{\mu})=0$. Then, $\gamma_{0}^{+}\equiv\left.\gamma^{+}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=(1-q)(\beta+\Delta)$ and $\gamma_{0}^{-}\equiv\left.\gamma^{-}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=q(\beta+\Delta)$. Note that when $q>\frac{1}{2}$, then $\gamma_{0}^{-}>\gamma_{0}^{+}$. Hence, if severe climate change is more likely a priori, then there are more distortions of good news than of bad news, and vice versa. Moreover, $\left.s^{+}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=\frac{\sigma^{2}\ln\left(1+\frac{\gamma}{q((\Delta+\beta)(1-q)-\gamma)}\right)}{2\mu}$ and $\left.s^{-}\right|_{\kappa(s,\mathbf{p},\tilde{\mu})=0}=\frac{\sigma^{2}\ln\left(\frac{\gamma(1-q)}{q(\Delta+\beta-\gamma)}\right)}{2\mu}.$ Hence, if $\gamma<\gamma_{0}^{+}$, then beliefs after $s>0$ are $q$ if $s\geq s^{+}_{0}$ and $\pi(s,\mathbf{p},\tilde{\mu})\in(q,1)$ else. Moreover, if $s<0$ and $\gamma<\gamma_{0}^{-}$, then beliefs are 00 if $s\in[s^{-}_{0},0)$ and $\pi(s)\in(0,q)$ else.
Note that all voters with $s<0$ hold beliefs weakly below Bayesian beliefs. Because of motivated beliefs, they vote to a greater extent informatively than they should, which is positive for information aggregation. Because $q\leq\tilde{\pi}$, they all vote informatively. If $\omega=0$, therefore a majority votes for policy 0. Hence, what matters for candidate incentives is the share of $s>0$ voters voting informatively. Because $q<\tilde{\pi}$, those with a belief of $q$ vote for policy 0. That is, all voters with $s>s_{0}^{+}$. Those with $s\in(0,s^{+})$ may vote either way.
The belief of a voter who chooses $\tilde{\mu}=\mu$ and receives signal $s=s^{+}_{0}$ is $\pi(s_{0}^{+},\mathbf{p},0)=\frac{\gamma+\Delta q+q}{\Delta+1}$. This is smaller than $\tilde{\pi}$ if
$\frac{\gamma+\Delta q+q}{\Delta+1}<\frac{1}{1+\beta}\Leftrightarrow\gamma<\tilde{\gamma}\equiv\frac{(\Delta+1)(1-q(1+\beta))}{\beta+1}>0.$
The belief of a voter who chooses $\tilde{\mu}=\mu$ and receives signal $s>0$ is
$\pi(s,\mathbf{p},\tilde{\mu})=\frac{q}{q+(1-q)e^{-\frac{2s\mu}{\sigma^{2}}}}.$
The signal at which such a voter starts to vote for policy 1 is the signal such that $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$. In particular, this signal realization is $s^{*}$, where $s^{*}\equiv\frac{\sigma^{2}\ln\left(\frac{1-q}{\beta q}\right)}{2\mu}$. If $\Phi\left(\frac{s-\mu}{\sigma}\right)$ is the CDF of the normal distribution with mean $\mu$ and standard deviation $\sigma$, the share of voters voting for policy 1 if $\omega=1$ is
$\mathcal{V}_{0}=\Phi\left(\frac{s^{+}_{0}-\mu}{\sigma}\right)-\Phi\left(\frac{s^{*}-\mu}{\sigma}\right).$
Recall that $s_{0}^{+}$ is increasing in $\gamma$, and hence $\mathcal{V}_{0}$ is monotonically increasing in $\gamma$. When $\gamma=0$, we know from our earlier analysis that voters all vote for policy 0 despite $s>0$. If $\gamma\rightarrow\infty$, no voters distorts $\mu$ and, by Assumption 2, a majority votes for policy 1 iff $\omega=1$. By continuity, there is a unique $\hat{\gamma}$ such that a majority votes for policy 0 if and only if $\gamma<\hat{\gamma}$. ∎
A.10 Proof of Lemma 4
@@
To see incentives, first consider how $W(s,\mathbf{p},\tilde{\mu})$ changes with $\tilde{\mu}$, when $\tilde{\mu}\geq\mu$. It is easy to show that $W(s,\mathbf{p},\tilde{\mu})$ is strictly concave in $\tilde{\mu}$ and thus an interior equilibrium exists for all $s$. The FOC for an interior optimum is
@@
$\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c(\tilde{\mu}-\mu).$
@@
$\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu).$
if $\tilde{\mu}<\mu$. In both derivatives, the first term’s sign is the opposite of the sign of $s$ (when $\Delta>1$). Hence, if $s>0$, then $\tilde{\mu}^{*}(s)<\mu$. Similarly, if $s<0$, we must have $\tilde{\mu}^{*}(s)>\mu$. If $s=0$, the first term is zero, and hence we need to have $\tilde{\mu}^{*}(s)=\mu$. Moreover, $\lim_{s\rightarrow\infty}\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu)$, and hence $\lim_{s\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$. Similarly, $\lim_{s\rightarrow-\infty}\left.\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-c(\tilde{\mu}-\mu)$, and hence also $\lim_{s\rightarrow-\infty}\tilde{\mu}^{*}(s)=\mu$.
@@
$-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}+\mu=\tilde{\mu}$
@@
$\Gamma=-\frac{2(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}$
@@
$\frac{\partial\Gamma}{\partial s}=\frac{-2(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left((1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)-qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)\right)}{c\sigma^{4}\left(q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+1\right)^{3}}>0,$
@@
$-\frac{2\tilde{\mu}^{3}(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}+\mu=\tilde{\mu}$
@@
$\Gamma^{\prime}=-\frac{2\tilde{\mu}^{3}(1-q)qs\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}$
@@
$\frac{\partial\Gamma^{\prime}}{\partial s}=\frac{2\tilde{\mu}^{3}(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)-(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)\right)}{c\sigma^{4}\left(\mu+\mu q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)\right)^{3}}>0$
@@
A.11 Proof of Proposition 6
The proof basically follows along the lines of the discussion in the text. We know from Lemma 4 that $\left|\tilde{\mu}(s)-\mu\right|$ decreases in $c$. Moreover, $\lim_{c\rightarrow\infty}\tilde{\mu}(s)=\mu$, because then both $\Gamma=0$ and $\Gamma^{\prime}=0$ (see the proof of Lemma 4). Moreover, as $c\searrow 0$, we are back in the scenario of Proposition 3. Finally note that both $\tilde{\mu}(s)$ and $\pi(s,\mathbf{p},\tilde{\mu})$ are continuous functions of both $s$ and $c$, and thus so is the share of voters voting for policy 1 in state $\omega=1$ if $p_{1}\neq p_{2}$. Hence, there must exist $\hat{c}$ such that the share of voters $\mathcal{V}_{0}$ voting for policy 1 in state 1 is just $\frac{1}{2}$. Moreover, $\text{Sign}\left[c-\hat{c}\right]=\text{Sign}\left[\mathcal{V}_{0}-\frac{1}{2}\right]$. This proves the proposition. ∎
Appendix B Additional Tables and Figures
Figure 5: Q-Q plots for the different regressions reported in Table 2.

### v5

@@ B.1 Cost from Imperfect Decision Making
$\begin{array}[]{rcl}W^{UC}(\tilde{\mu}=0)&=&-(1-\lambda)\left(b+\frac{(b+1)(q-1)}{q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1}\right)-\lambda q(b+\Delta)\\ W^{UC}(\tilde{\mu}=\mu_{c})&=&-\frac{\lambda(b+\Delta)}{b+1}\end{array}$
Note that
$\left.W^{UC}(\tilde{\mu}=0)\right|_{s=\tilde{s}}-W(\tilde{\mu}=\mu_{c})=\frac{\lambda(b+\Delta)(1-bq-q)}{b+1}>0$
for all $\lambda>0$. Moreover,
$\frac{\partial W^{UC}(\tilde{\mu}=0)}{\partial s}=-\frac{2(b+1)(1-\lambda)\mu(1-q)qe^{\frac{2\mu s}{\sigma^{2}}}}{\sigma^{2}\left[q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1\right]^{2}}<0.$
Therefore, if
$\begin{array}[]{c}\lim_{s\to\infty}W^{UC}(\tilde{\mu}=0)>W(\tilde{\mu}=\mu_{c})\Leftrightarrow b(\lambda+\lambda(-q)-1)-\Delta\lambda q>-\frac{\lambda(b+\Delta)}{b+1}\\ \Leftrightarrow\lambda>\tilde{\lambda}:=-\dfrac{b(b+1)}{\Delta(bq+q-1)+b(b(q-1)+q-2)}\in(0,1),\end{array}$
then any voter with $s>0$ will choose $\tilde{\mu}^{*}=0$. Conversely, if $\lambda<\tilde{\lambda}$, there exists $\hat{s}>\tilde{s}$, where
$\hat{s}=\frac{\sigma^{2}}{2\mu}\log\!\left(\frac{(q-1)\!\left[-\Delta\lambda+\lambda+b^{2}\lambda q+b\big((\Delta+1)\lambda q-1\big)+\Delta\lambda q-1\right]}{q\!\left[b+b^{2}(\lambda(q-1)+1)+b\lambda(\Delta q+q-2)+\Delta\lambda(q-1)\right]}\right),$
such that $\tilde{\mu}^{*}=0$ if $s\in[0,\hat{s}$] and $\tilde{\mu}^{*}=\mu_{c}$ if $s>\hat{s}$.
If $\lambda>\tilde{\lambda}$, then all results are as in the baseline version of the model. Otherwise, whether or not candidates will ignore their information depends on the complexity of climate change. For candidates to have incentives to offer policy 1, at least 50% of voters must receive $s>\hat{s}$ if $\omega=1$. If $\mu$ is sufficiently large, this condition is satisfied. However, if complexity is high, this will not hold, and the result from the baseline model continues to apply.
B.2 Cognitive Costs
Following the discussion in the main text, we now establish a useful lemma:
Lemma 2.
[…]
Proof.
@@ B.2 Cognitive Costs
To see incentives, first consider how $W^{CC}$ changes with $\tilde{\mu}$, when $\tilde{\mu}\geq\mu$. It is easy to show that $W^{CC}$ is strictly concave in $\tilde{\mu}$ and thus an interior equilibrium exists for all $s$. The FOC for an interior optimum is
@@
$\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c(\tilde{\mu}-\mu).$
@@
$\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu).$
if $\tilde{\mu}<\mu$. In both derivatives, the first term’s sign is the opposite of the sign of $s$ (when $\Delta>1$). Hence, if $s>0$, then $\tilde{\mu}^{*}(s)<\mu$. Similarly, if $s<0$, we must have $\tilde{\mu}^{*}(s)>\mu$. If $s=0$, the first term is zero, and hence we need to have $\tilde{\mu}^{*}(s)=\mu$. Moreover, $\lim_{s\rightarrow\infty}\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}<\mu}=-c\frac{\mu^{3}}{\tilde{\mu}^{3}}(\tilde{\mu}-\mu)$, and hence $\lim_{s\rightarrow\infty}\tilde{\mu}^{*}(s)=\mu$. Similarly, $\lim_{s\rightarrow-\infty}\left.\frac{\partial W^{CC}(s,\mathbf{p},\tilde{\mu})}{\partial\tilde{\mu}}\right|_{\tilde{\mu}\geq\mu}=-c(\tilde{\mu}-\mu)$, and hence also $\lim_{s\rightarrow-\infty}\tilde{\mu}^{*}(s)=\mu$.
@@
$\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}+\mu=\tilde{\mu}$
@@
$\Gamma=-\frac{2(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2{\tilde{\mu}}s}{\sigma^{2}}}}{\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}c}$
@@
$\frac{\partial\Gamma}{\partial s}=\frac{-2(1-q)q\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left((1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)-qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)\right)}{c\sigma^{4}\left(q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+1\right)^{3}}>0,$
@@
$\frac{2\tilde{\mu}^{3}(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}+\mu=\tilde{\mu}$
@@
$\Gamma^{\prime}=-\frac{2\tilde{\mu}^{3}(1-q)qs\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}}{c\mu^{3}\left(q\sigma\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)+\sigma\right)^{2}}$
@@
$\frac{\partial\Gamma^{\prime}}{\partial s}=\frac{2\tilde{\mu}^{3}(1-q)q\left[b+\Delta-(b+1)\kappa\right]e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(qe^{\frac{2\tilde{\mu}s}{\sigma^{2}}}\left(2\tilde{\mu}s-\sigma^{2}\right)-(1-q)\left(\sigma^{2}+2\tilde{\mu}s\right)\right)}{c\sigma^{4}\left(\mu+\mu q\left(e^{\frac{2\tilde{\mu}s}{\sigma^{2}}}-1\right)\right)^{3}}>0$
@@
The remaining discussion from the main text follows from the lemma.